OpenStudy (anonymous):

(squareroot of 2 - sqrt of 3 )^2

OpenStudy (xishem):

FOIL:\[(\sqrt{2}-\sqrt{3})^2 = (\sqrt{2}-\sqrt{3})(\sqrt{2}-\sqrt{3})\]\[2-\sqrt{2}\sqrt{3}-\sqrt{2}\sqrt{3}+3=5+2\sqrt{6}\]